1. Suppose that (X,Y) is uniformly distributed over the region {(x, y): 0 < \y< x < 1}. Find: a) the joint density of (X, Y);Need help with question 2 (not question 1)
Answer #1

1 2.) Lee S ={(44): 0 210)+14,41 :: Arca os S=4*3=2 (a) the joint density sum of (x,y) is $(14) = n-you Acom le 71745-1 71 D(c) clearly an annons fx (n) fy (0) + Fr, 121,8) = x,y are not independent. (&) Now B (²) Sn fx (m) dn (kan) du Se cun) de +